OpenStudy (studygurl14):

Can someone please help me? MEDAL! Evaluate the limit, if it exists. \(\huge\frac{6n^2+4n}{n^2+5}\) @ParthKohli @paki @confluxepic @Zale101 @EclipsedStar @AlexandervonHumboldt2 @iGreen @hartnn

Parth (parthkohli):

As it tends to infinity?

OpenStudy (studygurl14):

I don't know...

OpenStudy (studygurl14):

That's all the question says...

Parth (parthkohli):

Must be to infinity. In that case, divide both the numerator and denominator by n^2 and see how the thing would act as n approaches infinity

OpenStudy (anonymous):

n96254

Parth (parthkohli):

Actually, I'd prefer using L'Hopital's Rule.

OpenStudy (studygurl14):

So... \(\huge \frac{n^2(6+\frac{4}{n})}{n^2(1 + \frac{5}{n^2})}\) ???

OpenStudy (studygurl14):

like that?

OpenStudy (studygurl14):

Which would simplify to \(\huge \frac{6+\frac{4}{n}}{1+\frac{5}{n^2}}\)

Parth (parthkohli):

Or maybe you could do it the pure way.\[\lim \dfrac{6n^2 + 4n}{n^2 + 5} = \lim \dfrac{6n^2 + 4n}{n^2} = \lim 6 + 4/n\]

Parth (parthkohli):

And no, please ignore what I said earlier.

OpenStudy (studygurl14):

How did you get rid of the 5?

OpenStudy (anonymous):

There is 5..

Parth (parthkohli):

Actually, whatever, the other one works too. Since the denominator is approaching infinity either way, no one cares about the + 5.

OpenStudy (anonymous):

@StudyGurl14 just put n = infinity for what you have done..

OpenStudy (studygurl14):

So, I just say no limit? @ParthKohli ?

OpenStudy (anonymous):

There is..

Parth (parthkohli):

No. As n approaches infinity, what happens to your fraction?

OpenStudy (anonymous):

I am getting a finite value..

Parth (parthkohli):

What happens to 4/n and 5/n^2 as n starts to grow bigger and bigger?

OpenStudy (studygurl14):

The nubers get smaller

Parth (parthkohli):

Yeah, they get smaller. Ultimately, what number do they specifically reach (almost)?

OpenStudy (studygurl14):

0?

OpenStudy (anonymous):

you end up with: \[\large \frac{6 + \frac{4}{n}}{1 + \frac{5}{n^2}}\]

OpenStudy (anonymous):

yes, you are right..

OpenStudy (anonymous):

As you will put n = infinity, the fractions will become 0 and what are you left with now?

OpenStudy (studygurl14):

what do you mean "what are you left iwht now?"

OpenStudy (anonymous):

See, I try to show you :P

OpenStudy (anonymous):

\[\large \frac{6 + \frac{4}{n}}{1 + \frac{5}{n^2}} \implies \frac{6 + \frac{4}{\infty}}{1 + \frac{5}{\infty}} \implies \frac{6 + 0}{1 + 0}\]

OpenStudy (studygurl14):

Thank you @eta :)